# Mirror symmetry and rigid structures of generalized K3 surfaces

### v5

@@ 2.4. Moduli spaces and Torelli type theorem
Therefore the generalized K3 surfaces may be considered as geometric realizations of the points in the extended period domain $\widetilde{\mathfrak{D}}$. For this reason, we hereafter regard the generalized Calabi–Yau structures as elements in $H^{*}(M,\mathbb{C})$. The above Torelli type theorem will also serve as a foundation of our formulation of mirror symmetry in Section 4.
@@ 3.1. Lattices and K3 surfaces
In order to fix notation, let us begin with some basics of lattices and K3 surfaces. The hyperbolic lattice $U$ is the rank $2$ even unimodular lattice defined by the Gram matrix $\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}$. $E_{8}$ is the rank $8$ negative-definite even unimodular lattice defined by the corresponding Cartan matrix. We denote by $\Lambda_{K3}=U^{\oplus 3}\oplus E_{8}^{\oplus 2}$ the K3 lattice. For a lattice $L$, we write $L_{\mathbb{Q}}=L\otimes_{\mathbb{Z}}\mathbb{Q}$, and $L_{\mathbb{R}}$ and $L_{\mathbb{C}}$ are defined in a similar manner.
@@
$NS(S)=H^{1,1}(S,\mathbb{R})\cap H^{2}(S,\mathbb{Z}),\ \ \ T(S)=NS(S)^{\perp},$
where the orthogonal complement is taken in $H^{2}(S,\mathbb{Z})$. The Picard number is $\rho(S)=\mathrm{rank}(NS(S))$. The extended Néron-Severi lattice $NS^{\prime}(S)$ is the sublattice
$NS^{\prime}(S)=H^{0}(S,\mathbb{Z})\oplus NS(S)\oplus H^{4}(S,\mathbb{Z})\cong NS(S)\oplus U$
@@ 3.2. Néron–Severi lattice and transcendental lattice
$\widetilde{NS}(X)=\{\delta\in H^{*}(M,\mathbb{Z})\ |\ \langle\delta,\varphi^{\prime}\rangle=0\},$
$\widetilde{T}(X)=\{\delta\in H^{*}(M,\mathbb{Z})\ |\ \langle\delta,\varphi\rangle=0\}.$
It is important that we define $\widetilde{NS}(X)$ and $\widetilde{T}(X)$ on a completely equal footing, and our definition of $\widetilde{T}(X)$ differs from [6] (where it is defined as the orthogonal complement of $\widetilde{NS}(X)$). In our setting, the intersection $\widetilde{NS}(X)\cap\widetilde{T}(X)$ is in general non-trivial. Moreover, it is easy to see that
@@
In general, $H^{0}(M,\mathbb{Z}),H^{4}(M,\mathbb{Z})\not\subset\widetilde{NS}(X)$. This feature has an important consequence in our formulation of mirror symmetry (defined in Section 4).
@@
Let $\delta_{i}$ denote the degree $i$ part of $\delta\in H^{*}(M,\mathbb{Z})$.
@@
    If $\varphi^{\prime}=\sigma+B^{0,2}\wedge\sigma$, a twist of $\sigma$ by a $B$-field $B\in H^{2}(M,\mathbb{R})$, then
@@
    $=e^{B}NS^{\prime}(M)_{\mathbb{Q}}\cap H^{*}(M,\mathbb{Z}).$
    There is a natural inclusion $NS(M_{\sigma})\oplus H^{4}(M,\mathbb{Z})\subset\widetilde{NS}(X)$. The inclusion is strict if $B\in H^{2}(M,\mathbb{Q})$ and then $\widetilde{NS}(X)$ contains the lattice
@@
    In particular, if a $B$-field is absent, then
    $\widetilde{T}(X)=H^{2}(M,\mathbb{Z})_{\omega}\oplus\{\delta_{0}+\delta_{4}\ |\ \delta_{0}\int_{M}\omega^{2}=2\int_{M}\delta_{4}\},$
    where $H^{2}(M,\mathbb{Z})_{\omega}=\{\delta_{2}\ |\ \int_{M}\delta_{2}\wedge\omega=0\}$ is the space of $\omega$-primitive classes. Further, if $\omega^{2}\notin\mathbb{Q}$, we have $\widetilde{T}(X)=H^{2}(M,\mathbb{Z})_{\omega}$.
@@ 3.3. Complex rigidity
The singular K3 surfaces are also known as complex rigid K3 surfaces in the sense that they admit no complex deformation keeping the maximal Picard number $20$ (or equivalently, the rank of the transcendental lattice minimal $2$). We first give a generalization of complex rigid K3 surfaces.
@@
    $M_{\sigma}$ is a complex rigid K3 surface,
@@
In other words, $X$ is a rational $B$-field $B^{\prime}$ shift of a complex rigid K3 surface equipped with a complexified Kähler structure.
@@
Let $X=(\varphi,\varphi^{\prime})$ be complex rigid. By the classification of hyperKähler structures (comments below Remark 2.7), $\varphi$ must be of type $A$. Hence we can write $X=e^{B^{\prime}}(\lambda e^{B+\sqrt{-1}\omega},\sigma)$, where $Y=(\lambda e^{B+\sqrt{-1}\omega},\sigma)$ is a generalized K3 surface, i.e. $B$ is a closed real $(1,1)$-form and either ±$\pm\omega$ is a hyperKähler form for $\sigma$. On the other hand, for $\varphi^{\prime}=\sigma+B^{\prime}\wedge\sigma$, $\mathrm{rank}(L_{\varphi^{\prime}})=2$ if and only if $\mathrm{rank}(T(M_{\sigma}))=2$ and $B^{\prime}$ is rational. This can been seen as follows. If $\mathrm{rank}(L_{\varphi^{\prime}})=2$, considering the degree $2$ part, we must have $\mathrm{rank}(L_{\sigma})=2$. Then we may write $\sigma=\alpha v_{1}+\sqrt{-1}\beta v_{2}$ for $\alpha,\beta\in\mathbb{R}$ and $v_{1},v_{2}\in L_{\sigma}=T(M_{\sigma})$. Then
$\varphi^{\prime}=\alpha(v_{1}+B^{\prime}\wedge v_{1})+\sqrt{-1}\beta(v_{2}+B^{\prime}\wedge v_{2})$
and in order for $\mathrm{rank}(L_{\varphi^{\prime}})=2$, the $B$-field $B^{\prime}$ must be rational. The converse is also true. ∎
@@ 3.4. Kähler rigidity
In light of mirror symmetry, a rigid Kähler structure has been anticipated for a long time. A Kähler rigidity comparable with a complex rigidity naturally appears in the framework of generalized Calabi–Yau structures.
@@
$v_{1}=(1,0,-n),\ \ v_{2}=(0,H,0)\in NS^{\prime}(S).$
Then
@@
$=v_{1}+\sqrt{-1}v_{2}\in(\mathbb{Z}v_{1}+\mathbb{Z}v_{2})_{\mathbb{C}}\subsetneq NS^{\prime}(S)_{\mathbb{C}}$
On the other hand, for $\epsilon^{2}\notin\mathbb{Q}$,
@@
Hence there is no proper sublattice $L\subsetneq NS^{\prime}(S)$ such that $e^{\sqrt{-1}\epsilon H}\in L_{\mathbb{C}}$. Therefore the Kähler structure $H$ is not continuously deformable in such a way that $\mathrm{rank}(L_{e^{\sqrt{-1}\epsilon H}})=2$. This calculation illustrates that a generalized Calabi–Yau structure $e^{B+\sqrt{-1}\omega}$ is able to detect subtle integral structures of the Kähler moduli space. In other words, by non-linearize a Kähler structure $B+\sqrt{-1}\omega$ by $e^{B+\sqrt{-1}\omega}$, we can make a new interesting lattice $L_{e^{B+\sqrt{-1}\omega}}$ visible.
@@
Again, the condition $\mathrm{rank}(\widetilde{T}(X))=22$ is equivalent to $\mathrm{rank}(L_{\varphi})=2$. We will next provide a characterization of the Kähler rigidity. In contrast to the complex rigidity, the hyperKähler structure $\varphi^{\prime}$ can be either of type $A$ or $B$.
@@
First, $B$ needs to be rational, and hence so is $\omega^{2}$. Then we may write the symplectic class $\omega=\kappa H$ for $\kappa^{2}\in\mathbb{Q}$ and $H\in H^{2}(M,\mathbb{Z})$ with $H^{2}>0$. Indeed, in this case, there exist $m,n\in\mathbb{N}$ such that
@@ 4.1. Mirror symmetry for K3 surfaces
We begin with a brief review of mirror symmetry for lattice polarized K3 surfaces developed by Dolgachev in [3].
@@
For an integer $1\leq\rho\leq 19$, let $K\subset\Lambda_{K3}$ be sublattice of signature $(1,\rho-1)$. A $K$-polarized K3 surface is a pair $(S,i)$ of a K3 surface $S$ and a primitive lattice embedding $i:K\hookrightarrow NS(S)$ such that $i(K)$ contains an ample class.
@@
For a sublattice $K\subset\Lambda_{K3}$ of signature $(1,\rho-1)$, assume there exists a lattice $L$ such that $K^{\perp}=L\oplus U$. Then the family $\mathcal{S}$ of $K$-polarized K3 surfaces and the family $\mathcal{S}^{\vee}$ of $L$-polarized K3 surfaces are mirror symmetric.
In the above setting, $K$ is of signature $(1,19-\rho)$, and for a generic $K$-polarized K3 surface $S$ and a generic $L$-polarized K3 surface $S^{\vee}$, we have
@@
They are considered as a duality between the algebraic cycles of $S$ and the transcendental cycles of $S^{\vee}$ as lattices, and vice versa (duality of Yukawa couplings). Mirror symmetry can also be investigated on the level of the moduli spaces.
The crucial assumption in Dolgachev’s formulation is the existence of a decomposition of the form $K^{\perp}=L\oplus U$, which is a reflection of $NS^{\prime}(S)=NS(S)\oplus U$ under mirror symmetry. Hence the assumption is unremovable as long as we consider geometry where the point class and fundamental class play special roles.
@@
It is also worth noting that there are geometric explanations for such a decomposition; the hyperbolic lattice $U$ corresponds to
@@ 4.2. Problems
Although Dolgachev’s formulation works very nicely in many cases, it has some problems and cannot be a definitive formulation. For example, $NS^{\prime}(S)$ and $T(S)$ are not really symmetric; the minimal rank of $NS^{\prime}(S)$ is $3$ (as long as $S$ is algebraic) while the minimal rank of $T(S)$ is $2$. Moreover, the assumption $K^{\perp}=L\oplus U$ does not hold in general. There are two types of problems.
@@
    There are decompositions of the form $K^{\perp}=L\oplus U(m)$ but for $m\geq 2$. Then the duality between the algebraic cycles of a $K$-polarized K3 surface and the transcendental cycles of an $L$-polarized K3 surface fails. Nevertheless, the duality still holds with $\mathbb{Q}$-coefficients and the moduli correspondence works without problems [3].
@@
    The situation is worse if $K^{\perp}$ is of signature $(2,0)$ and hence cannot be of the form $U(m)$ for any $m$. Such a $K$ is given by $K=NS(S)$ for a singular K3 surface $S$. The natural question “How can we understand mirror symmetry for singular K3 surfaces?” has been a mysterious problem for a long time. In fact, for a singular K3 surface, the Kähler moduli space is of dimension 20 and complex moduli space is of dimension 0, and there seems no mirror partner as a K3 surface.
@@
We will provide answers to these problems in the framework of generalized K3 surfaces in Sections 4.5 and 4.6.
Remark 4.4.
In [2], Aspinwall–Morrison identified the moduli space of certain non-linear $\sigma$-models on a K3 surface with $\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R}))\cong\mathrm{O}(4,20)/(\mathrm{SO}(2)\times\mathrm{SO}(2)\times\mathrm{O}(20))$. Geometrically, to a K3 surface $M_{\sigma}$ with a Kähler class $\omega$, we associate an orthogonal pair of positive oriented 2-planes
$(\langle\mathrm{Re}(\sigma),\mathrm{Im}(\sigma)\rangle_{\mathbb{R}},\langle\mathrm{Re}(e^{\sqrt{-1}\omega}),\mathrm{Im}(e^{\sqrt{-1}\omega})\rangle_{\mathbb{R}})\in\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R})).$
However, there are “non-geometric points” that are not obtained in this way. Mirror symmetry, which is an involution of $\mathrm{Gr}_{2,2}^{po}(H^{*}(M,\mathbb{R}))$, does not preserve the geometric points. Therefore it is known that the mirror partner of a K3 surface is not necessarily a K3 surface.

### v6

@@ 2.4. Moduli spaces and period domains
Therefore the generalized Calabi–Yau structures of hyperKähler type may be considered as geometric realizations of the points in the extended period domain $\widetilde{\mathfrak{D}}$. For this reason, we hereafter regard a generalized Calabi–Yau structure of hyperKähler type as an element in $H^{*}(M,\mathbb{C})$.
@@ 3.1. Lattices and K3 surfaces
In order to fix notation, let us begin with some basics of lattices and K3 surfaces. The hyperbolic lattice $U$ is the rank $2$ even unimodular lattice defined by the Gram matrix $\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}$. $E_{8}$ is the rank $8$ negative-definite even unimodular lattice defined by the corresponding Cartan matrix. We denote by $\Lambda_{K3}:=U^{\oplus 3}\oplus E_{8}^{\oplus 2}$ the K3 lattice. For a lattice $L$, we write $L_{\mathbb{Q}}:=L\otimes_{\mathbb{Z}}\mathbb{Q}$, and $L_{\mathbb{R}}$ and $L_{\mathbb{C}}$ are defined in a similar manner.
@@
$NS(S):=H^{1,1}(S,\mathbb{R})\cap H^{2}(S,\mathbb{Z}),\ \ \ T(S):=NS(S)^{\perp},$
where the orthogonal complement is taken in $H^{2}(S,\mathbb{Z})$. The Picard number is $\rho(S):=\mathrm{rank}(NS(S))$. The extended Néron-Severi lattice $NS^{\prime}(S)$ is the sublattice
$NS^{\prime}(S):$ $=H^{0}(S,\mathbb{Z})\oplus NS(S)\oplus H^{4}(S,\mathbb{Z})$
$\cong NS(S)\oplus U$
@@ 3.2. Néron–Severi lattice and transcendental lattice
$\widetilde{NS}(X):=\{\delta\in H^{*}(M,\mathbb{Z})\ |\ \langle\delta,\varphi^{\prime}\rangle=0\},$
$\widetilde{T}(X):=\{\delta\in H^{*}(M,\mathbb{Z})\ |\ \langle\delta,\varphi\rangle=0\}.$
In [13] the Picard group $\mathrm{Pic}(\varphi)$ and the transcendental lattice $T(\varphi)$ are defined for a generalized Calabi–Yau structure $\varphi$. Our definition of the transcendental lattice is different from [13], where it is defined as the orthogonal complement of the Picard group.
It is important that we define $\widetilde{NS}(X)$ and $\widetilde{T}(X)$ on a completely equal footing, and hence the intersection $\widetilde{NS}(X)\cap\widetilde{T}(X)$ may be non-trivial. Moreover, it is easy to see that
@@
For example, $\varphi$ is chosen generically in $\widetilde{NS}(X)_{\mathbb{C}}$, then $\widetilde{NS}(X)^{\perp}=\widetilde{T}(X)$.
@@
In general, $H^{0}(M,\mathbb{Z}),H^{4}(M,\mathbb{Z})\not\subset\widetilde{NS}(X)$. This feature has an important consequence in our formulation of mirror symmetry (Section 4).
The following examples will play an important role in Section 4. (cf. Example 4.4 of [13] )
@@
We denote by $\delta_{i}$ the degree $i$ part of $\delta\in H^{*}(M,\mathbb{Z})$.
@@
    If $\varphi^{\prime}=\sigma+B^{0,2}\wedge\sigma$, a $B$-field shift of $\sigma$ by $B\in H^{2}(M,\mathbb{R})$, then
@@
    $=e^{B}NS^{\prime}(M_{\sigma})_{\mathbb{R}}\cap H^{*}(M,\mathbb{Z}).$
    There is a natural inclusion $NS(M_{\sigma})\oplus H^{4}(M,\mathbb{Z})\subset\widetilde{NS}(X)$. The inclusion is strict if $B\in H^{2}(M,\mathbb{Q})$, and then $\widetilde{NS}(X)$ contains the lattice
@@
    In particular, if $\sigma$ is a holomorphic 2-form for a complex structure on $M$, then for a generic choice of $B,\omega\in NS(M_{\sigma})$
    $\widetilde{T}(X)=T(M_{\sigma}).$
@@ 3.3. Complex rigidity
The singular K3 surfaces, also known as complex rigid K3 surfaces, do not admit any complex deformation while maintaining the maximal Picard number $20$ (or equivalently, the minimal rank $2$ of the transcendental lattice). We first provide a generalization of the singular K3 surfaces.
@@
    $M_{\sigma}$ is a singular K3 surface,
@@
In other words, $X$ is a rational $B$-field shift of a singular K3 surface equipped with a complexified Kähler structure.
@@
Let $X=(\varphi,\varphi^{\prime})$ be complex rigid. By the classification of hyperKähler structures (comments below Remark 2.8), $\varphi$ must be of type $A$. Hence we can write $X=e^{B^{\prime}}(\lambda e^{B+\sqrt{-1}\omega},\sigma)$, where $Y=(\lambda e^{B+\sqrt{-1}\omega},\sigma)$ is a generalized K3 surface, i.e. $B$ is a closed real $(1,1)$-form and either ±$\pm\omega$ is a hyperKähler form for $\sigma$. On the other hand, for $\varphi^{\prime}=\sigma+B^{\prime}\wedge\sigma$, $\mathrm{rank}(L_{\varphi^{\prime}})=2$ if and only if $\mathrm{rank}(T(M_{\sigma}))=2$ and $B^{\prime}$ is rational. This can be seen as follows. If $\mathrm{rank}(L_{\varphi^{\prime}})=2$, there exist $a,b\in\mathbb{R}$ and $u_{1},u_{2}\in H^{*}(M,\mathbb{Z})$ such that
$\varphi^{\prime}=\sigma+B^{\prime}\wedge\sigma=au_{1}+\sqrt{-1}bu_{2}.$
For $i=1,2$, let $v_{i}=(u_{i})_{2}\in H^{2}(M,\mathbb{Z})$ be the degree 2 part of $u_{i}$. Then we may write $\sigma=av_{1}+\sqrt{-1}bv_{2}$ and hence $\mathrm{rank}(L_{\sigma})=2$, which means $\mathrm{rank}(T(M_{\sigma}))=2$. Moreover, we have
(1) $\varphi^{\prime}=e^{B}\sigma=a(v_{1}+B^{\prime}\wedge v_{1})+\sqrt{-1}b(v_{2}+B^{\prime}\wedge v_{2}),$
and in order for $\mathrm{rank}(L_{\varphi^{\prime}})=2$, the $B$-field $B^{\prime}$ must be rational. Conversely, assume $\mathrm{rank}(T(M_{\sigma}))=2$ and $B^{\prime}$ is rational, then there exist $a,b\in\mathbb{R}$ and $v_{1},v_{2}\in T(M_{\sigma})$ such that $\sigma=av_{1}+\sqrt{-1}bv_{2}$, and $\varphi^{\prime}$ is of the form of the equation (1). Since $B^{\prime}$ is rational, $v_{1}+B^{\prime}\wedge v_{1},v_{2}+B^{\prime}\wedge v_{2}\in H^{*}(M,\mathbb{Q})$ and hence $\mathrm{rank}(L_{\varphi^{\prime}})=2$. ∎
@@ 3.4. Kähler rigidity
A Kähler rigidity comparable with a complex rigidity naturally appears in the framework of generalized Calabi–Yau structures.
@@
$v_{1}:=(1,0,-n),\ \ v_{2}:=(0,H,0)\in NS^{\prime}(S),$
where the notation follows the decomposition $NS^{\prime}(S)=H^{0}(S,\mathbb{Z})\oplus NS(S)\oplus H^{4}(S,\mathbb{Z})$. Then
@@
$=v_{1}+\sqrt{-1}v_{2}\in(\mathbb{Z}v_{1}+\mathbb{Z}v_{2})_{\mathbb{C}}\subsetneq NS^{\prime}(S)_{\mathbb{C}}.$
Here
$\mathbb{Z}v_{1}+\mathbb{Z}v_{2}\cong\langle 2n\rangle^{\oplus 2},$
where $\langle k\rangle$ denotes the lattice of rank $1$ generated by $v$ with $v^{2}=k$. On the other hand, for $\epsilon^{2}\notin\mathbb{Q}$,
@@
Hence there is no proper sublattice $L\subsetneq NS^{\prime}(S)$ such that $e^{\sqrt{-1}\epsilon H}\in L_{\mathbb{C}}$. Therefore the Kähler structure $H$ is not continuously deformable in such a way that $\mathrm{rank}(L_{e^{\sqrt{-1}\epsilon H}})=2$. This new sublattice $L_{e^{B+\sqrt{-1}\omega}}$ is comparable with the transcendental lattice, which is the minimal sublattice $L\subset H^{2}(M_{\sigma},\mathbb{Z})$ such that $\sigma\in L_{\mathbb{C}}$.
@@
Again, the condition $\mathrm{rank}(\widetilde{T}(X))=22$ is equivalent to $\mathrm{rank}(L_{\varphi})=2$. We will next provide a characterization of the Kähler rigidity. Recall that, in contrast to the complex rigidity, the hyperKähler structure $\varphi^{\prime}$ can be either of type $A$ or $B$.
@@
First, $B$ needs to be rational, and hence so is $\omega^{2}$. Then we may write the symplectic class $\omega=\kappa H$ for $\kappa^{2}\in\mathbb{Q}$ and $H\in H^{2}(M,\mathbb{Z})$ with $H^{2}>0$. In fact, in this case, there exist $m,n\in\mathbb{N}$ such that
@@ 4.1. Mirror symmetry for K3 surfaces
We begin with a brief summary of mirror symmetry for lattice polarized K3 surfaces originally introduced by Dolgachev in [7]. Fix an integer $1\leq\rho\leq 19$.
@@
Let $K$ be a lattice of signature $(1,\rho-1)$. A $K$-polarized K3 surface is a pair $(S,i)$ of a K3 surface $S$ and a primitive lattice embedding $i:K\hookrightarrow NS(S)$ such that $i(K)$ contains an ample class.
@@
For a lattice $K$ of signature $(1,\rho-1)$, assume there exists a lattice $L$ such that
$K^{\perp}=L\oplus U.$
Then the family $\mathcal{S}$ of $K$-polarized K3 surfaces and the family $\mathcal{S}^{\vee}$ of $L$-polarized K3 surfaces are mirror symmetric.
In the above setting, $L$ is of signature $(1,19-\rho)$ and uniquely determined with $L^{\perp}=K\oplus U$ (independent of the chose of a decomposition and an embedding) [7]. Then, for a generic $K$-polarized K3 surface $S$ and a generic $L$-polarized K3 surface $S^{\vee}$, we have
@@
They are considered as a duality between the algebraic cycles of $S$ and the transcendental cycles of $S^{\vee}$ as lattices, and vice versa (duality of Yukawa couplings). Mirror symmetry can also be explored at the level of moduli spaces. See for example [7, 8, 11].
The crucial assumption in Dolgachev’s formulation is the existence of a decomposition of the form $K^{\perp}=L\oplus U$, which reflects the decomposition
$NS^{\prime}(S)=NS(S)\oplus\big{(}H^{0}(S,\mathbb{Z})\oplus H^{4}(S,\mathbb{Z})\big{)}$
under mirror symmetry. Hence the assumption is indispensable as long as we consider geometry where the point class and fundamental class play special roles.
@@
It is also worth noting that there are geometric explanations for such a decomposition; the hyperbolic lattice $U$ corresponds to:
@@
Therefore, a choice of such $U$ corresponds to a choice of a mirror K3 surface.
@@ 4.2. Problems
Although Dolgachev’s formulation works well in many cases, it has a few issues and cannot be considered definitive. First of all, $NS^{\prime}(S)$ and $T(S)$ are not truly symmetric; the minimal rank of $NS^{\prime}(S)$ is $3$ (as long as $S$ is algebraic) while the minimal rank of $T(S)$ is $2$. Additionally, the assumption $K^{\perp}=L\oplus U$ does not hold in general. These issues fall into two categories.
@@
    There are decompositions of the form $K^{\perp}=L\oplus U(m)$ but for $m\geq 2$. Examples of such K3 surfaces are constructed for examples in [10]. The duality between the algebraic cycles of a $K$-polarized K3 surface and the transcendental cycles of an $L$-polarized K3 surface holds only over $\mathbb{Q}$.
@@
    The situation is worse if $K^{\perp}$ is of signature $(2,0)$ and hence cannot be of the form $U(m)$ for any $m\geq 1$. Such a lattice $K$ is given by $K=NS(S)$ for a singular K3 surface $S$. Therefore, the natural question “How can we understand mirror symmetry for singular K3 surfaces?” has been a mysterious problem for a long time. Indeed, for a singular K3 surface, the Kähler moduli space is of dimension 20 and complex moduli space is of dimension 0, and there seems no mirror partner as a K3 surface.
@@
We will settle these problems in the framework of generalized K3 surfaces in Sections 4.5 and 4.6.
More generally, the existence of rigid Calabi–Yau manifolds poses a serious challenge to mirror symmetry. Various approaches have been proposed to tackle this issue, but there seem currently no universally agreed-upon formulation. See for example [4, 6, 16, 18].